OpenStudy (anonymous):

160x/-420 =48x/720 What do I do from here?

OpenStudy (anonymous):

is this \[\frac{160x}{-420}\]?

OpenStudy (anonymous):

yes and 48x/720

OpenStudy (anonymous):

-8x/21 = x/15 -8x/(3*7) = x/(5*3) -40x = 7x -47x = 0 x = 0

OpenStudy (anonymous):

160x/-420=48x/720

OpenStudy (nottim):

CoolSector did it right I think, dizzy. Look closely; he simplified your fractions.

OpenStudy (nottim):

Actually, I have no clue. What's the full question?